# Algebraic geometry of bubbling Kähler metrics

### v1

@@ 2.2.3. Negative deformation
We discuss the notion of negative weight deformation in the sense of [CH22] and equivalently rephrase more algebraically in terms of the generalized test configurations of [Od24b]. This finishes the proof of the nontrivial direction of [Od24b], and answers the Open Question 1.4 (2) of [CH22]. Note that, in the context of the previous section, the degenerations $\mathcal{X}^{\prime}_{0}\rightsquigarrow W$ and $\mathcal{X}^{\prime\prime}_{0}\rightsquigarrow C$ are both negative weight deformations by the construction.
@@
    ([Od24b, Lemma 2.19]) Take an arbitrary algebraic negative weight deformation (cf., Definition 2.6 of [Od24b]) of a Fano cone $T\curvearrowright(Y,\xi)$ as $p\colon\mathcal{Y}\twoheadrightarrow U_{\tau}$, with $\xi\in-\tau$. Then, the general fiber of $p$ is a negative $\xi$-weight deformation in the original sense of Definition 1.7 of [CH22].
@@ 2.3. Analogue for Kähler-Ricci solitons - Properness of moduli and bubbling solitons
The latter is (essentially) constructed in [BLXZ23, §6, Theorem 6.1 (and 6.4)], as loc.cit Theorem 6.1 proves the existence of $\mathcal{M}^{\rm Fano}_{n,V,(1,)h^{+}}$ which is our $\mathcal{M}_{n,V,h^{+}}$, and it follows from their Theorem 6.4 that $\mathcal{M}_{n,V,h}$ is a locally closed substack of $\mathcal{M}_{n,V,h^{+}}$. Indeed, by the same theorem, it even follows that $\mathcal{M}_{n,V,h^{+}}$ has a finite stratification by of $\mathcal{M}_{n,V,h^{\prime}}$ for finite possibilities of $h^{\prime}(\geq h)$. Now we hope that the following analogue of [Che24, 1.1], [Od24a, Theorem 3.5] hold.
@@
Also, the moduli stack $\mathcal{M}^{s}_{n,V,T,\xi}$ of (modified) K-semistable pair $(X\curvearrowleft T,\xi)$ exists. If $\mathbbm{k}=\mathbb{C}$ and the moduli stack $\mathcal{M}^{s}_{n,V,T,\xi}$ admits a coarse good moduli space $M^{s}_{n,V,T,\xi}$, its complex analytification $M^{s}_{n,V,T,\xi}(\mathbb{C})$ is compact.
@@
In conclusion, we can expect very similar methods of constructing the coarse (good) moduli space $\mathcal{M}^{s}_{n,V,T,\xi}$ of (weak) Kähler-Ricci solitons, to the construction of the coarse good moduli space of the moduli stack of the Calabi-Yau cones in [Od24a]. Also note that the above algebraic compactness result is compatible with the precompactness type result by [PSS23, GPSS18].
2.4. Re-construction of minimal bubble via coordinates
Now, finishing the digression in the previous subsection, we come back to the original context of affine klt varieties and the local bubbling problem. From what follows, for simplicity, we work over $\mathbbm{k}=\mathbb{C}$ under the Setups 1 (a) $\land$ 2 $\land$ 3. Further we assume that $\mathcal{X}\to\Delta$ is a restriction of algebraic morphism to a unit disk $\Delta$. As it is often obvious, many of the following discussions also work verbatim for algebraic families over other base fields $\mathbbm{k}$ (as in Setup 1 (b)).
In this subsection, we discuss a priori different (but conjecturally equivalent) and more practical reconstruction of the minimum bubble, in terms of coordinates and defining equations. For that, we work again under the setup 1, 2, and further suppose that $X$ and $W$ is obtained inside the affine space $\mathbb{C}^{l}$ and $v_{X}$ is obtained by weights $(w_{1},\cdots,w_{l})\in\mathbb{R}_{>0}^{l}$ corresponding to the coordinates $z_{i}$ of $\mathbb{C}^{l}$. For certain sufficiently divisible $N$ and $w_{i}$s, this is always achievable ([DS17, §3.3]). We would like to put the weight on $t$, the base coordinate, as $w_{0}\in\mathbb{R}_{>0}$ and $w:=(w_{0},\cdots,w_{l})$. In this subsection, we further assume $\mathcal{X}$ is the restriction of affine morphism over $\mathbb{A}_{t}^{1}$ for technical simplicity.
@@ 2.4. Re-construction of minimal bubble via coordinates
We now temporarily fix the extended embedding of $\mathcal{X}$ into $\Delta\times\mathbb{C}^{l}$, and consider $\mathcal{W}:=\mathcal{W}(w_{0};w_{1},\cdots,w_{l}):=\mathop{\mathrm{Spec}}\nolimits(gr_{w_{0},w_{1},\cdots,w_{l}}(\mathcal{O}_{\mathcal{X},x}))$ as a family over $\mathbb{A}^{1}_{t}(\mathbb{C})=\mathbb{C}$ where $t$ is the coordinate of $\Delta$, as in [dBS23]. Here, $gr_{w_{0},w_{1},\cdots,w_{l}}(\mathcal{O}_{\mathcal{X},x})$ is the natural graded ring
$\oplus_{m\geq 0}\{\overline{f}\in\mathcal{O}_{\mathcal{X},x}\mid f\in\mathcal{O}_{\Delta\times\mathbb{C}^{l},0},w(f)\geq m\}/\{\overline{f}\in\mathcal{O}_{\mathcal{X},x}\mid f\in\mathcal{O}_{\Delta\times\mathbb{C}^{l},0},w(f)>m\},$
where for $f=\sum a_{i_{0},i_{1},\cdots,i_{N}}t^{i_{0}}\prod_{1\leq j\leq N}z_{j}^{i_{j}}$, the weight $w(f)$ of $f$ means $\min\{\sum_{j}w_{j}i_{j}\mid a_{i_{0},i_{1},\cdots,i_{N}}\neq 0\}$ and $\overline{f}$ means the restriction.
@@
    there is a critical value $w_{\rm cri}\in\mathbb{R}_{>0}$, such that
@@
This proposition is inspired by [Sun23, dBS23] and the discussions with the authors.
@@
As we assumed in the beginning of this subsection, $\mathcal{X}\subset\Delta\times\mathbb{C}^{l}$ is cut out by some defining polynomials $F_{1},\cdots,F_{m}$ of $z_{1},\cdots,z_{N},t$. We (again) take enough $F_{i}$s to form a universal Gröbner basis of the generating ideal ([BCRV22, §5.1]). As $x=0\in\mathcal{X}$, these do not contain constant terms. Then $\mathcal{W}\subset\mathbb{A}_{t}^{1}\times\mathbb{C}^{l}$ is cut out by the initial terms ${\rm in}_{w_{0},w_{1},\cdots,w_{l}}(F_{i})$ of $F_{i}$ with respect to the weights $w_{0},w_{1},\cdots,w_{l}$. Thus in particular, it admits the natural $\mathbb{C}^{*}$-action which equivariantly projects down to $\mathbb{A}_{t}^{1}$ with the weight $w_{0}$.
@@
Conjecture 2.24 (cf., [Sun23, dBS23]).
@@
For a good choice of $N$, embedding of $\mathcal{X}$ to $\Delta\times\mathbb{C}^{l}$ and $w_{1},\cdots,w_{l}$, $\mathcal{W}(w_{\rm cri};w_{1},\cdots,w_{l})\to\mathbb{A}^{1}$ is a test configuration $\mathcal{T}$, which degenerates the central fiber $\mathcal{X}^{\prime}_{\rm min,0}$ of $\mathcal{X}$ in Theorem 2.4 isotrivially to $W$ (compare the proof of Theorem 2.4).
@@
Note that [dBS23] implicitly confirmed the above conjecture in classical examples in low dimension which inspire above (we also did some similar calculations to loc.cit, confirming the above in higher dimensional examples e.g. generic hypersurfaces). Indeed, to obtain the minimum bubble in this general, we need a good choice of extension of the embedding $X$ into $\mathbb{C}^{l}$ to $\mathcal{X}$ to $\Delta\times\mathbb{C}^{l}$ as the construction of [Sun23, Lemma 4.6] and the following bad example suggests.
Example 2.25.
Consider the case $N=3$, when $\mathcal{X}$ is defined by $((z_{1}+tz_{3})^{2}+(z_{2}+tz_{3})^{2}=z_{3}^{m}+t^{h})$ in $\Delta\times\mathbb{C}^{l}$ for $h\gg m\geq 2$. Then if we still consider the coordinates $z_{i}$s and $t$,
  • $\mathcal{W}(w_{0};w_{1},\cdots,w_{l})\to\mathbb{A}^{1}$ for $w_{0}>w_{\rm cri}$ is $(z_{1}^{2}+z_{2}^{2}=z_{3}^{m})\simeq W\times\mathbb{A}^{1}$,
  • $\mathcal{W}(w_{\rm cri};w_{1},\cdots,w_{l})\to\mathbb{A}^{1}$ is $((z_{1}+tz_{3})^{2}+(z_{2}+tz_{3})^{2}=z_{3}^{m})\simeq W\times\mathbb{A}^{1}$.
2.5. Re-construction of minimal bubble via valuation

### v2

@@ 2.2.5. Negative weight deformation as scaling down
We discuss the notion of negative weight deformation in the sense of [CH22] and equivalently rephrase more algebraically in terms of the generalized test configurations of [Od24b]. In op.cit, we also called it scale down deformation. As the name suggests, we aim to provide algebro-geometric formulation for the opposite analogue of Theorem 2.3. This finishes the proof of the nontrivial direction of [Od24b], and answers the Open Question 1.4 (2) of [CH22]. Note that, in the context of the previous section, the degenerations $\mathcal{X}^{\prime}_{0}\rightsquigarrow W$ and $\mathcal{X}^{\prime\prime}_{0}\rightsquigarrow C$ are both negative weight deformations by the construction.
@@
    ([Od24b, Lemma 2.19]) Take an arbitrary algebraic negative weight deformation (cf., Definition 2.6 of [Od24b]) of a Fano cone $T\curvearrowright(Y,\xi)$ as $p\colon\mathcal{Y}\twoheadrightarrow U_{\tau}$, with $\xi\in\tau$. Then, the general fiber of $p$ is a negative $\xi$-weight deformation in the original sense of Definition 1.7 of [CH22].
@@ 2.3. Analogue for Kähler-Ricci solitons - Properness of moduli and bubbling solitons
The latter is (essentially) constructed in [BLXZ23, §6, Theorem 6.1 (and 6.4)], as op.cit Theorem 6.1 proves the existence of $\mathcal{M}^{\rm Fano}_{n,V,(1,)h^{+}}$ which is our $\mathcal{M}_{n,V,h^{+}}$, and it follows from their Theorem 6.4 that $\mathcal{M}_{n,V,h}$ is a locally closed substack of $\mathcal{M}_{n,V,h^{+}}$. Indeed, by the same theorem, it even follows that $\mathcal{M}_{n,V,h^{+}}$ has a finite stratification by of $\mathcal{M}_{n,V,h^{\prime}}$ for finite possibilities of $h^{\prime}(\geq h)$. Now we hope that the following analogue of [Che24, 1.1], [Od24a, Theorem 3.5] hold.
@@
Also, the moduli stack $\mathcal{M}^{s}_{n,V,T,\xi}$ of (modified) K-semistable pair $(X\curvearrowleft T,\xi)$ exists. If $\mathbbm{k}=\mathbb{C}$ and the moduli stack $\mathcal{M}^{s}_{n,V,T,\xi}$ admits a categorical good  [^16] We refer to [Stacksproject, 106.12] (resp., [Alp13]) for the definition of categorical moduli (resp., good moduli). Before, the author has often used the term “coarse moduli” for the exactly same meaning as the categorical moduli in older papers, i.e., in a broader sense than e.g., [Mum65], not assuming the condition of bijectivity of geometric points as in op.cit. moduli space $M^{s}_{n,V,T,\xi}$, its complex analytification $M^{s}_{n,V,T,\xi}(\mathbb{C})$ is compact.
@@
By the above compactness theorem, we can expect very similar methods of constructing the good moduli space $\mathcal{M}^{s}_{n,V,T,\xi}$ of (weak) Kähler-Ricci solitons, to the construction of the good moduli space of the moduli stack of the Calabi-Yau cones in [Od24a]. Also note that the above algebraic compactness result is compatible with the precompactness type result by [PSS23, GPSS18].
2.4. Construction of minimal bubble via coordinates
Now, finishing the digression in the previous subsection, we come back to the original context of affine klt varieties and the local bubbling problem.
In this subsection, we discuss a priori different (but conjecturally equivalent) and more practical reconstruction of the minimum bubble, in terms of coordinates and defining equations. From what follows, for notational simplicity, we work over the Setup 1 (b) $\land$ 2 $\land$ 3. It is almost verbatim, by setting $R:=\mathbb{C}[[t]]^{\rm conv}$, how to discuss a complex analytic version for the Setup 1 (a) $\land$ 2 $\land$ 3.
As in [Od24a, Od24b], we assume (without loss of generality) $X$ and $W$ are embedded into the affine space $\mathbb{A}_{\mathbbm{k}}^{l}$ and the isotrivial degeneration $X\rightsquigarrow W$ is obtained by weights $(w_{1},\cdots,w_{l})\in\mathbb{R}_{>0}^{l}$ corresponding to the coordinates $z_{i}$ of $\mathbb{A}_{\mathbbm{k}}^{l}$. We would like to also put a weight on $t$, the base coordinate, as $w_{0}\in\mathbb{R}_{>0}$.
We further fix the extended embedding of $\mathcal{X}$ into $\Delta\times\mathbb{A}^{l}$ which maps $\sigma$ to the 00-section. By the arguments of proof of Theorem 2.4 and Appendix, this setup is not restrictive. Now, we consider
$\mathcal{W}:=\mathcal{W}(w_{0};w_{1},\cdots,w_{l}):=\mathop{\mathrm{Spec}}\nolimits(gr_{w_{0},w_{1},\cdots,w_{l}}(\mathcal{O}_{\mathcal{X},x})),$
@@ 2.4. Construction of minimal bubble via coordinates
as a family over $\mathbb{A}^{1}_{t}$ where $t$ is the coordinate of $\Delta$, as in [dBS24]. Here, $gr_{w_{0},w_{1},\cdots,w_{l}}(\mathcal{O}_{\mathcal{X},x})$ is the natural graded ring
(11) $\oplus_{m\geq 0}\{\overline{f}\in\mathcal{O}_{\mathcal{X},x}\mid f\in\mathcal{O}_{\Delta\times\mathbb{A}^{l},0},w(f)\geq m\}/\{\overline{f}\in\mathcal{O}_{\mathcal{X},x}\mid f\in\mathcal{O}_{\Delta\times\mathbb{A}^{l},0},w(f)>m\},$
where for $f=\sum a_{i_{0},i_{1},\cdots,i_{N}}t^{i_{0}}\prod_{1\leq j\leq N}z_{j}^{i_{j}}$, the weight $w(f)$ of $f$ means $\min\{\sum_{j}w_{j}i_{j}\mid a_{i_{0},i_{1},\cdots,i_{N}}\neq 0\}$ and $\overline{f}$ means the restriction. Note that this construction (11) is a variant of the graded ring discussed in e.g., [LXZ22, XZ22] but its finite typeness over $\mathbbm{k}$ is easy in this setup. Indeed, if the ideal sheaf for $\mathcal{X}\subset\mathbb{A}^{k}\times\Delta$ is $I$, and $w(\mathcal{O}_{\Delta\times\mathbb{A}^{k},0})$ consists of $m_{1}<m_{2}<\cdots$, if we set $\{f\in\mathcal{O}_{\Delta\times\mathbb{A}^{l},0}\mid f\in\mathcal{O}_{\Delta\times\mathbb{A}^{l},0},w(f)\geq m_{i}\}=:R_{i}$, we have
(12) $\{\overline{f}\in\mathcal{O}_{\mathcal{X},x}\mid f\in\mathcal{O}_{\Delta\times\mathbb{A}^{l},0},w(f)\geq m_{i}\}/\{\overline{f}\in\mathcal{O}_{\mathcal{X},x}\mid f\in\mathcal{O}_{\Delta\times\mathbb{A}^{l},0},w(f)\geq m_{i+1}\}$
(13) $=R_{i}/(R_{i+1}+\{f\in\mathcal{O}_{\Delta\times\mathbb{A}^{l},0}\mid f\in\mathcal{O}_{\Delta\times\mathbb{A}^{l},0},w(f)\geq m_{i}\}\cap I)\twoheadleftarrow R_{i}/R_{i+1},$
for each $i=1,2,\cdots$ so that ${\rm gr}_{w_{0},\cdots,w_{l}}(\mathcal{O}_{\mathbb{A}^{k},0})=\oplus_{i}R_{i}/R_{i+1}\simeq\mathbbm{k}[z_{1},\cdots,z_{l},t]$ is a finite type algebra over $\mathbbm{k}$.
@@
    there is a (natural) critical value $w_{\rm cri}\in\mathbb{R}_{>0}$, such that
@@
This proposition is inspired by [Sun23, dBS24] and the discussions with the authors.
@@
As we assumed in the beginning of this subsection, $\mathcal{X}\subset\Delta\times\mathbb{A}^{l}$ is cut out by some defining polynomials $F_{1},\cdots,F_{m}$ of $z_{1},\cdots,z_{N},t$. We (again) take enough $F_{i}$s to form a universal Gröbner basis of the generating ideal ([BCRV22, §5.1]). As $x=0\in\mathcal{X}$, these do not contain constant terms. Then $\mathcal{W}\subset\mathbb{A}_{t}^{1}\times\mathbb{A}^{l}$ is cut out by the initial terms ${\rm in}_{w_{0},w_{1},\cdots,w_{l}}(F_{i})$ of $F_{i}$ with respect to the weights $w_{0},w_{1},\cdots,w_{l}$. Thus in particular, it admits the natural $\mathbb{G}_{m}$-action which equivariantly projects down to $\mathbb{A}_{t}^{1}$ with the weight $w_{0}$.
@@
We now prepare a key observation to connect the previous deformation theoretic construction of bubbling (Theorem 2.4, 2.6) to approaches via coordinates (or valuations).
In particular, we analyze a general structure of $T$-equivariant morphism $\widetilde{\rm ST}_{\mathbbm{k}[[t]]}(\tau^{\prime})\to M$ (see [Od24b, §3] and the previous section for the notation) which appears in the proof of Theorem 2.4. Here, $\tau^{\prime}$ is a rational polyhedral cone of $N\oplus\mathbb{Z}$ which projects as PL isomorphism (in particular, bijectively) to $\tau\subset N_{\mathbb{R}}$. The rational polyhedral cone $\tau^{\prime}$ includes a vector $(w_{0},\xi)\in(\mathbb{Z}\oplus N)\otimes\mathbb{R}$, which projects to $\xi\in\tau\subset N\otimes\mathbb{R}$.
In the context (proof) of Theorem 2.4, we were concerned with particular $w_{0}$ which is naturally determined by the other data including $\xi$. Nevertheless, here we consider arbitrary $w_{0}$ and corresponding $T$-equivariant morphism $\widetilde{\rm ST}_{\mathbbm{k}[[t]]}(\tau^{\prime})\to M$ and analyze the obtained degenerating families from different perspectives.
Lemma 2.25 (Relation with weighted blow ups).
For $T\curvearrowright M$ and $\Delta\to M$ in the proof of Theorem 2.4, where the latter specifies $\mathcal{X}\to\Delta$ and its (00-)section $\sigma$, we consider an arbitrary $T(=N\otimes\mathbb{G}_{m})$-equivariant morphism $\overline{f}\colon\widetilde{\rm ST}_{\mathbbm{k}[[t]]}(\tau^{\prime})\to M$ for some $\tau^{\prime}$ as the previous two paragraphs right above here. We take a primitive integral $\xi^{\prime}=(\xi,w_{0})\in(\tau^{\prime}\cap N\oplus\mathbb{Z})$ (thus approximating the original direction $\xi^{\prime}$ of the arguments in [Od24b, 3.8] with rational rank $r$, in general) which projects to integral $\xi\in\tau\cap N$.
  1. ($\mathrm{i}$)
    We consider the rank $1$ algebraic subtorus of $T\times\mathbb{G}_{m}$ (resp., the rank $1$ algebraic subtorus of $T$) associated to $\mathbb{R}\xi^{\prime}\cap(N\oplus\mathbb{Z})$ (resp., $\mathbb{R}\xi\cap N$) as $(\mathbb{G}_{m}\simeq)T_{\xi^{\prime}}\subset T\times\mathbb{G}_{m}$ (resp., $(\mathbb{G}_{m}\simeq)T_{\xi}\subset T)$). Consider the corresponding affine family $\tilde{\mathcal{X}}_{\overline{f},\tau^{\prime}}\to\widetilde{\rm ST}_{\mathbbm{k}[[t]]}(\tau^{\prime})$ to $\overline{f}$, restrict it to the closure of $T_{\xi^{\prime}}$-orbit in $\widetilde{\rm ST}_{\mathbbm{k}[[t]]}(\tau^{\prime})$ and denote as $\mathcal{X}^{\prime}(w_{0},\tau^{\prime})\to\mathbb{A}^{1}$.
    Then, if $w_{0}=1$, $\mathcal{X}^{\prime}(w_{0}=1,\tau^{\prime})$ is the weighted blow up of $\mathcal{X}$ for weights $(w_{0},w_{1},\cdots,w_{l})$. For general $w_{0}\geq 1$, $\mathcal{X}^{\prime}(w_{0},\tau^{\prime})$ is the normalization of the base change of the weighted blow up of $\mathcal{X}$ for weights $(w_{0},w_{1},\cdots,w_{l})$ with respect to the $w_{0}$-th power map $\Delta\xrightarrow{s\mapsto s^{w_{0}}=t}\Delta$. In particular, for large enough $w_{0}$, the central fiber (over 00) of $\mathcal{X}^{\prime}(w_{0},\tau^{\prime})\to\mathbb{A}^{1}$ does not depend on $w_{0}$.
  2. ($\mathrm{ii}$)
[…]
By proof of Theorem 2.4, $\mathcal{X}_{\rm min}^{\prime}$ and $\mathcal{W}$ obtained therein can be regarded as particular cases of the above construction ($\mathrm{i}$), ($\mathrm{ii}$).
proof of Lemma 2.25.
Firstly, note that $\tilde{\mathcal{X}}_{\overline{f},\tau^{\prime}}\to\widetilde{\rm ST}_{\mathbbm{k}[[t]]}(\tau^{\prime})$ is affine and faithfully-flat finite type from the definition. We first prove ($\mathrm{i}$). Take the weights of $T_{\xi}$ on the coordinates $z_{1},\cdots,z_{l}$ of $\mathbb{A}^{l}$ as $w_{1},\cdots,w_{l}$. Then, for a fixed $c\in\mathbbm{k}^{*}$, $\mathcal{X}^{\prime}(w_{0},\tau^{\prime})$ is isomorphic to $\overline{\{(t^{-w_{0}},\cdots,t^{-w_{l}})\cdot\mathcal{X}_{t^{w_{0}}c}\}}_{t\in(\Delta\setminus 0)}$ as closed subschemes of $\mathbb{A}^{l}\times_{\mathbbm{k}}\Delta=\mathbb{A}_{\mathbbm{k}[[t]]}^{l}$, where $\cdot$ means the $T_{\xi}$-action, $0\in\Delta$ means the closed point (the origin) and the overline means the (Zariski) closure inside $\mathbb{A}_{z_{1},\cdots,z_{l}}^{l}\times_{\mathbbm{k}}\Delta$. Note that presence of $T_{\xi}$-action on $\mathbb{A}_{\mathbbm{k}}^{l}$ makes the above construction independent of $c$. If $w_{0}=1$, equivalently, it ($\mathcal{X}^{\prime}(w_{0}=1,\tau^{\prime})$) is the restriction of the Zariski closure of $\mathcal{X}\setminus X$ in $\mathbb{A}^{l}_{\frac{z_{1}}{t^{w_{1}}},\cdots,\frac{z_{l}}{t^{w_{l}}}}$ (note the different coordinates), and equivalently, it is nothing but the weighted blow up of $\mathcal{X}$ with weights $(w_{0}=1,w_{1},\cdots,w_{l})$. Hence the assertion of ($\mathrm{i}$) holds in this case. For general $w_{0}$ case, let us denote the $w_{0}$-th power map of the base $\Delta_{s}\xrightarrow{s\mapsto s^{w_{0}}=t}\Delta_{t}$, which we use for base change. Then the normalization of ${\rm Bl}_{(w_{0},\cdots,w_{l})}(\mathbb{A}_{\mathbbm{k}[[t]]}^{l})\times_{\mathbbm{k}[[t]]}\mathbbm{k}[[s]]$ is ${\rm Bl}_{(1,w_{1},\cdots,w_{l})}(\mathbb{A}_{\mathbbm{k}[[s]]}^{l})$ as it is easy to see ring-theoretically. Therefore, as closed their subschemes, there is a finite morphism from ${\rm Bl}_{(1,w_{1},\cdots,w_{l})}(\mathcal{X}\times_{\mathbbm{k}[[t]]}\mathbbm{k}[[s]])$ to ${\rm Bl}_{(w_{0},w_{1},\cdots,w_{l})}(\mathcal{X})\times_{\mathbbm{k}[[t]]}\mathbbm{k}[[s]]$ which is automatically the normalization. Indeed, note that the former is normal because of the previous arguments for $w_{0}=1$ case and the fact that it is normal as the total space of a fiberwise normal family by our assumption on $M$. Hence, the assertion of ($\mathrm{i}$) follows from the case of $w_{0}=1$ and we conclude its proof, except for the last statements, which now follows from Proposition 2.24.
Next, we prove ($\mathrm{ii}$). When we consider small enough perturbation of $(w_{0},w_{1},\cdots,w_{l})$ which takes values in $\mathbb{Q}\oplus N_{\mathbb{Q}}$, note that $\lim_{t\to 0}(t^{-dw_{0}},\cdots,t^{-dw_{l}})\cdot X_{t^{dw_{0}}c}$ for $d\in\mathbb{Z}_{>0}$ such that $dw_{i}\in\mathbb{Z}$ (in the sense of Zariski closure inside $\mathbb{A}^{l}_{\mathbbm{k}}\times_{\mathbbm{k}}\Delta=\mathbb{A}_{\mathbbm{k}[[t]]}^{l}$) remain the same. This fact simply follows from the definition of $\tilde{\mathcal{X}}_{\overline{f},\tau^{\prime}}\to\widetilde{\rm ST}_{\mathbbm{k}[[t]]}(\tau^{\prime})$ as what corresponds to $\overline{f}$. On the other hand, by our definition of graded ring in (11) (compare analogous [LX18, 2.10]) and our arguments in the proof of Proposition 2.24, ${\rm gr}_{w_{0},w_{1},\cdots,w_{l}}(\mathcal{O}_{\mathcal{X},x})$ also does not change as a ring (ignoring the grading) for a certain perturbation to $w_{0},w_{1},\cdots,w_{l}\in\mathbb{Q}\oplus N_{\mathbb{Q}}$. Indeed, it also follows from the description of ${\rm gr}_{w_{0},w_{1},\cdots,w_{l}}(\mathcal{O}_{\mathcal{X},x})$ in terms of initial ideals (see [Eis13, Chapter 15]). So we can and do assume $w_{i}$ are all integers. Then note that for any $c\in\mathbbm{k}$, admitting $c=0$, $\lim_{t\to 0}(t^{w_{0}},\cdots,t^{w_{l}})\cdot X_{c\cdot t^{w_{0}}}\subset\mathcal{W}_{c}$ for any $c$ by the proof of the above ($\mathrm{i}$). Here, the above limit is as close subscheme of $\mathbb{A}^{l}$. Therefore,
$\lim_{t\to 0}(t^{w_{0}},\cdots,t^{w_{l}})\cdot\mathcal{X}=\mathcal{W}(w_{0},\cdots,w_{l})\times_{\mathbb{A}^{1}_{\mathbbm{k}}}\mathbbm{k}[[t]],$
as closed subschemes of $\mathbb{A}_{\mathbbm{k}[[t]]}^{l}$, since the left hand side is contained in the right hand side while both are irreducible hence they are the same. As they are both restrictions of affine test configurations to over $\mathbbm{k}[[t]]$, one can extend the above isomorphism via the horizontal $\mathbb{G}_{m}$-action to $\mathcal{W}\simeq{\rm Spec}({\rm gr}_{\xi^{\prime}}(\mathcal{O}_{\mathcal{X},x}))$.
Now, we apply the above Lemma 2.25 to the situation of (proof of) Theorem 2.4.
Corollary 2.26 (cf., Conjectures in [dBS24, Sun23]).
@@
In the Setup 1, for a good choice of embedding of $\mathcal{X}$ to $\Delta\times\mathbb{A}_{\mathbbm{k}}^{l}$ and $w_{1},\cdots,w_{l}$, $\mathcal{W}(w_{\rm cri};w_{1},\cdots,w_{l})\to\mathbb{A}^{1}$ (of Proposition 2.24) is a test configuration $\mathcal{T}$, which degenerates the central fiber $\mathcal{X}^{\prime}_{\rm min,0}$ of $\mathcal{X}$ in Theorem 2.4 isotrivially to $W$ (compare the proof of Theorem 2.4).
proof of Corollary 2.26.
Firstly, consider the valuation $v_{X}$ of $\mathcal{O}_{X,x}$ which minimizes $\widehat{\rm vol}(-)$. It exists by [Blu18] and is quasi-monomial [Xu20] and further has a finitely generated graded ring ${\rm gr}_{v_{X}}(\mathcal{O}_{X,x})$ [XZ22], as it matches to the differential geometric perspectives provided in [DS17]. Therefore, as explained in [Od24a, §2], one can take homogeneous generators of ${\rm gr}_{v_{X}}(\mathcal{O}_{X,x})$ and lifts to global sections of $\mathcal{O}_{\mathcal{X}}$ which embeds $\mathcal{X}$ into $\mathbb{A}^{1}\times\mathbb{A}^{l}$. Then, there are weights $w_{1},\cdots,w_{l}$ on the fiber-direction coordinates on $\mathbb{A}^{l}$ and the according construction in the proof of Theorem 2.4 in particular gives $w_{0}$ so that corresponding family $\mathcal{W}(w_{0},\cdots,w_{l})$ is the (nontrivial) negative weight deformation in Theorem 2.4 ($\mathrm{iv}$) with the general fiber $\mathcal{X}^{\prime}_{\rm min,0}$. Further, from the construction of proof of Theorem 2.4, it follows that for any $0<w<w_{0}$, we have $\mathcal{W}(w,w_{1},\cdots,w_{l})$ is $W\times\mathbb{A}^{1}$. On the other hand, by Lemma 2.25 ($\mathrm{ii}$), this $\mathcal{W}(w,\cdots,w_{l})$ can be understood as ${\rm gr}_{w,\cdots,w_{l}}(\mathcal{O}_{\mathcal{X},x})$ in the sense of (11) in general. Therefore, combining with Proposition 2.24, we obtain that $w_{0}=w_{\rm cri}$ so that Corollary 2.26 follows. ∎
@@
The above corollary allows explicit descriptions of the K-semistable bubbling $\mathcal{X}^{\prime}_{\rm min,0}$ in terms of weighted blow ups, and is inspired by discussions with deBorbon-Spotti (cf., [dBS24, Remark 11]), S.Sun and S.Donaldson. The author appreciates them. Note also that [dBS24] implicitly confirmed the above conjecture in classical examples in low dimension which inspire above (we also did some similar calculations to loc.cit, confirming the above in higher dimensional examples e.g. generic hypersurfaces). Indeed, in order to obtain the differential minimum bubble in general, we need a good choice of extension of the embedding $X$ into $\mathbb{C}^{l}$ to $\mathcal{X}$ to $\Delta\times\mathbb{C}^{l}$ as the construction of [Sun23, Lemma 4.6] suggests.
2.5. Valuative approach